OpenStudy (anonymous):

Which is an equation in point-slope form of the line that passes through the points (7, 5) and (-4, -1)? A. y + 1 = 6/7(x + 4) B. y + 1 = 6/11(x + 4) C. y + 4 = 11/6(x + 1) D. y - 1 = 6/11(x - 4)

OpenStudy (anonymous):

\[y-y _{1}/x-x _{1}\]

OpenStudy (anonymous):

so say subtracting the second from the first we have 5-(-1) / 7-(-4) the y values over the x values

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

slope is rise over run or y over x right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so a slope of 1/3 would be up 1 on the y axis and right 3 which is the x axis

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so to find the slope of two points you use the equation y-y1/ x-x1 to get a y / x value

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you choose one of the point pairs and subtract those from the other pairs

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so if we decide to subtract the second pair from the first pair then the first set of (x,y) values are (7,5) and the second set are (-4,-1) so you would subtract the second y value from the first and the second x value from the first

OpenStudy (anonymous):

you go either way it doesnt matter if you choose to subtract the first values from the second as long as you remain consistant

OpenStudy (anonymous):

but subtracting the second y value from the first you have 5-(-1) for the rise and for the run you subtract the second x value from the first so 7-(-4) for the run

OpenStudy (anonymous):

this gives you a slope of 5-(-1) / 7-(-4) = 6/11 right?

OpenStudy (anonymous):

are you following so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so with a slope of 6/11 to find the equation you insert one of the set of points and the slope you just found into the formula y-y1 = m(x+x1)

OpenStudy (anonymous):

m is the slope and y1 and x1 are the values you will replace so choosing a set of points say the second set and inserting those into the formula we have y-(-1)=6/11(x-(-4))

OpenStudy (anonymous):

you cannot use the y value from set 2 and the x value from set 1

OpenStudy (anonymous):

can you close the problem please